OpenStudy (anonymous):

FAN AND MEDAL!! what is the fourth term of the expansion of the binomial (2x+5)to the fifth power?

Directrix (directrix):

That would be C(5,3) * (2x)^(5-3) * (5)^3 = ?

Directrix (directrix):

C(5,3) = 10 C(5,3) * (2x)^(5-3) * (5)^3 = 10* (2x)^3 * 5^3 = ? @timbra

Directrix (directrix):

Here's a link that leads to a review of this: http://www.purplemath.com/modules/binomial2.htm

Directrix (directrix):

C(5,3) = 10 C(5,3) * (2x)^(5-3) * (5)^3 = 10* (2x)^2 * 5^3 = ? @timbra

OpenStudy (anonymous):

thanks!!!!

Directrix (directrix):

What did you get for the final answer when you simplified this?